Circular No. 5832

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET WEST-KOHOUTEK-IKEMURA (1993o) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery

of this comet with the Spacewatch telescope on Kitt Peak. On July 20 the observations through three airmasses and in twilight showed a coma of diameter 8′′. The indicated correction to the prediction on MPC 16381 (ephemeris on MPC 21961) is ΔT = +0.04 day.

1993 UT  α_{2000}  δ_{2000}  m_{1}
July 20.47149  3^{h}09^{m}56.^{s}15  −14^{o}05′40′′.3  20.1
20.47495  3 09 56.57  −14 05 38.2  20.1
21.46759  3 11 45.98  −14 02 36.0  20.1
21.47095  3 11 46.18  −14 02 34.5
21.47414  3 11 46.58  −14 02 34.7

SUPERNOVA 1993J IN NGC 3031 R. Pressberger and H. M. Maitzen, Institute for Astronomy, Vienna,

report further 6-color photometry (including Strömgren, given as stated on cf. IAUC 5822): June 29.93 UT, u = +2.1, v = +1.61, b = +1.55, g_{1} = +1.35, g_{2} = +1.19, y = +1.06; July 17.93, +2.2, +1.80, +1.81, +1.67, +1.41, +1.44. These data, combined with their previous ones, reveal a linear decrease in brightness as of May 12 with the following rates (± 0.0005 magnitude/day): y, 0.0182 mag/day; g_{2}, 0.0134; g_{1}, 0.0160; b, 0.0138; v, 0.0078. The u brightness increased from May 12 to June 1 and then remained practically constant until July 17.

A. W. Neely, Silver City, NM, reports that he has revised his preliminary unfiltered CCD magnitudes of SN 1993J so that they correspond to Johnson V magnitudes, based on subsequent observations of standard fields with the same CCD and Johnson V, Kron-Cousins R, and RG9 glass I filters: Mar. 28.30 UT, V = 13.45 ± 0.15; 29.15, 11.05 ± 0.2; 30.05, V = 10.68 ± 0.2, R = 10.68 ± 0.2, I = 10.49 ± 0.2; Apr. 1.11, 11.10 ± 0.03, 10.85 ± 0.04, 10.72 ± 0.04; 2.10, 11.47, 11.20, 10.88; 19.20, 10.83, 10.51, 10.52; 20.20, 10.83, 10.50, 10.49; 21.20, 10.90, 10.52, 10.47; May 1.17, 11.80, 11.16, 10.88; 21.15, 12.47, 11.78, 11.36; June 11.17, 12.87, 12.23, 11.79; 22.16 12.93, 12.45, 12.03; 23.16, 12.98, 12.53, 12.06; 29.16, 13.08 ± 0.05, 12.59 ± 0.06, 12.04 ± 0.06; July 8.16, 13.30, 12.81, 12.37.

1993 July 23 Daniel W. E. Green